OpenStudy (anonymous):

for each of these equations: i put the equation into the form (x-h)^2 + (y-k)^2 = r^2 ii state the centre and the radius a) x^2+6x+y^2+2y+=6 (b) x^2-8X+y^2+4y=29 c) x^2+2x+y^2+14y=14 (d)x^2-4x+y^2+6y=12 e) x^2+10x+y^2-12y=3 (f) x^2-8x+y^2-2y=19

OpenStudy (dumbcow):

you have to complete the square... here is general example \[x^{2} +bx = 4\] \[x^{2} +bx+(\frac{b}{2})^{2} = 4 +(\frac{b}{2})^{2}\] \[(x +\frac{b}{2})^{2} = 4+(\frac{b}{2})^{2}\]